OpenStudy (anonymous):

What is the mean absolute deviation for the data? 30, 30, 35, 35, 40 A. 3.2 B. 5 C. 10 D. 34

OpenStudy (anonymous):

HELEPP ME MHELP ME

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

find the mean (30+30+35+35+40) / 5 = 170/5 = 34 now take the absolute value of the difference from each set of data | 30 - 34 | = | -4 | = 4 |30 - 34| = | -4| = 4 | 35 - 34| = 1 | 35 - 34| = 1 | 40 - 34| = 6 now take the mean (average) of these numbers (4 + 4 + 1 + 1 + 6) / 5 16/5 = 3.2 <=== this is you MAD (mean absolute deviation)

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

i just need the answer HELLO

OpenStudy (anonymous):

T-T

OpenStudy (anonymous):

answer is 3.2